# Artificial Intelligence HCI Human-Computer Interaction

### v2

@@ 5.3. High AI Involvement: AI-Led Workflows
At this level, AI takes the lead in the QDA process, performing tasks autonomously with varying degrees of human oversight.
@@
AI primarily conducts QDA tasks, while human researchers provide input, refine outputs, or make adjustments at key stages. As P8 noted, “The AI suggests a codebook, and I give feedback that the tool uses to make an updated suggestion.” This role allows AI to handle much of the workload while human researchers retain control over decisions. Cohn et al. (Cohn et al., 2024) propose a human-in-the-loop prompt engineering approach to collaborative discourse analysis. Dai et al. (Dai et al., 2023) develop an LLM-in-the-loop framework to develop an initial codebook.
AI with Human-in-the-Loop [Uncaptioned image] Figure 12. AI takes the lead in QDA, e. g., creating a codebook draft (left) or performing initial coding (right), with human oversight throughout. :
@@
AI autonomously performs QDA tasks, but human approval is required before finalizing outputs. This ensures AI contributions are vetted for quality and integrity. P13 suggested, “AI could apply my codebook, and I manually review a random subsample.” Recent studies use LLMs as additional coders that are evaluated in a final step by computing agreement metrics with human coders (Tai et al., 2024; Bano et al., 2024; Xiao et al., 2023). Drapal et al. (Drápal et al., 2023) propose a framework in which an LLM predicts and applies code and themes with intermediate reviews by the researcher.
AI with Human Approval [Uncaptioned image] Figure 13. AI creates a codebook draft (top) or assigns codes (bottom). A researcher approves before decisions are finalized and changes are applied. :
@@
AI with Conditional Autonomy [Uncaptioned image] Figure 14. AI autonomously handles routine coding but alerts researchers when it encounters data requiring nuanced understanding or deviating from the specified criteria. :
Full AI Delegation
At the highest level of AI involvement, AI autonomously manages the entire QDA process without human intervention. OpenAI’s Viable is an AI model that can be used for analysis of unstructured data at scale, for example, in businesses that require regular customer feedback evaluations. Rasheed et al. (Rasheed et al., 2024) propose a multi-agent system that takes data sources, autonomously performs the chosen approach of QDA, and outputs a complete analysis. AI tools for QDA must be highly customizable to adapt to the varied tasks and needs of researchers in qualitative research. One promising avenue for achieving this flexibility is through the concept of agentic AI, which envisions AI systems as autonomous agents capable of performing tasks, making decisions, and interacting with users in a dynamic and customizable manner  (Shavit et al., 2023). While this level of autonomy may be suitable for large-scale data processing, it raises concerns around data validity, interpretive depth, and the loss of human oversight (Jiang et al., 2021). Participants stressed the importance of maintaining agency and control over AI-driven processes.
5.4. Summary: AI Tools for QDA Should Balance Support with Ethical and Practical Considerations
While HCI researchers were open to exploring the potential of AI support for QDA, they emphasized the need for responsible design that respects research ethics, quality, and researcher control. On the basis of our interview results and prior literature, we synthesized a framework of varying AI involvement – ranging from minimal to high – that allows researchers to choose the appropriate level of AI assistance based on their specific needs and project context. This approach ensures that AI enhances productivity and provides valuable insights without undermining the researcher’s agency. By incorporating AI in a way that is transparent, ethically sound, and aligned with established research practices, researchers can responsibly integrate AI into QDA workflows, ensuring it serves as a complement rather than a replacement for human expertise.
@@ 6. Discussion
6.1. Synthesizing QDA Practices in CSCW
Our findings extend prior discussions of AI-assisted qualitative research by grounding them in a systematic account of how CSCW and HCI researchers actually conduct QDA in practice.
[…]
While AI-supported QDA offers novel opportunities, a major concern regarding the quality of outputs and inherent biases prevails. This is particularly important when analyzing non-English data (Heseltine and Clemm von Hohenberg, 2024; Suter and Meckel, 2024), engaging minority groups, or conducting research in the Global South, as AI is not “culturally agnostic” (Paulus and Marone, 2024) and often privileges dominant or high-frequency representations over others (Feuston and Brubaker, 2021). Perspectives from the Global South and majority world are often marginalized in HCI and CSCW (Kumar and Karusala, 2021; Sturm et al., 2015; Linxen et al., 2021)and in related disciplines (Hasegawa et al., 2024; Sturm et al., 2015). AI-supported QDA risks amplifying this imbalance by embedding dominant interpretative frameworks and overlooking locally relevant meanings (Kumar and Karusala, 2021), particularly when models are trained on Global North or WEIRD datasets (Linxen et al., 2021). These risks are compounded by ongoing challenges in HCI and CSCW around citational justice (Kumar and Karusala, 2021) and transparency of research practices (Wacharamanotham et al., 2020). Addressing them is essential for building culturally sensitive approaches to AI-supported QDA.
6.2. Putting Our Framework Into Practice
How our framework might be used
Both prior research and our empirical findings highlight the dual need to respect the interpretive, collaborative nature of QDA while exploring how AI might responsibly augment it. Building on these insights, we propose a nuanced framework for human-AI collaboration in QDA, with varying roles for AI across increasing levels of involvement, tailored to the specific needs of researchers performing QDA. We derive tangible scenarios for AI support that go beyond coding to include tasks such as data preprocessing, onboarding new team members, mediating conflicts, and acting as a conversational partner. We present this framework as flexible rather than prescriptive, recognizing that many factors must be considered when integrating AI into qualitative research. Our findings contribute to ongoing CSCW discussions around responsible, human-centered AI design. They offer timely insights into how HCI researchers are beginning to re-evaluate the role of AI in light of rapidly advancing LLM capabilities.
Our framework aims to empower those who want to incorporate AI tools into QDA, but are unsure of how to approach it or where to intervene. A naive approach is to jump towards AI doing the qualitative coding itself, but this was a contentious topic for our participants. Instead, our framework offers numerous ways AI tools can be incorporated in ways that facilitate QDA without compromising research autonomy. Researchers who are curious about incorporating AI tools such as LLMs but are worried about how this may taint the QDA process can strive for minimal AI involvement, such as AI as a training tool. For companies developing QDA software (e. g., nVivo or MAXQDA), many of which are already adding AI tools to their software, our framework hopes to inspire additional insights into the numerous ways AI tools can be developed to support QDA. Determining the appropriate level and role of AI integration requires careful consideration of factors like research goals, data types, and QDA methods. Below, we outline key considerations for integrating AI into QDA.
Determining the appropriate level of AI involvement
When designing AI-assisted QDA workflows, one should consider what AI tools are capable of and what they should do (Feuston and Brubaker, 2021; Lubars and Tan, 2019). For example, large-scale or structured datasets may benefit from AI’s ability to handle pattern recognition, while nuanced or sensitive data might necessitate limited AI involvement to maintain interpretive depth and ethical considerations. Research goals – whether they focus on exploration, theory generation, or confirmation – also dictate how AI can best support QDA. While in early-stage research, AI can support the identification of patterns and themes, in later stages, AI could be useful for consistency checks or validating human-coded data.
Deciding when AI should lead or support
Related work shows that coding performance can depend on various factors of complexity, e. g., the required level of judgment and interpretation  (Ortloff et al., 2023; Kirsten et al., 2024). While higher task complexity is an indicator of reduced AI involvement, routine and repetitive tasks like transcription offer the potential for AI leadership, freeing researchers to focus on more complex interpretive tasks. AI should be employed where its capabilities align with the complexity of the task. For example, when the model shows high certainty in a task, it might take a more leading role, while in tasks involving ambiguity or high interpretive judgment, AI could step back. In addition, researchers’ comfort and familiarity with AI tools should guide their level of involvement (Lubars and Tan, 2019; Gauthier and Wallace, 2022). Recent work (Bhatt and Sargeant, 2024; Gauthier and Wallace, 2022; Lam et al., 2024) discusses how restricting AI outputs or displaying performance disclaimers might aid the effective use of AI.
Identifying stages where human oversight is essential
Critical review points must be established to ensure that human oversight is maintained where necessary. Depending on the level of AI involvement, researchers should determine at which stages oversight is crucial -— such as after each coding pass, document review, or when noteworthy patterns are identified. In cases where AI operates autonomously in a conditional mode, scenarios that require human intervention need to be defined, e. g., under user-specified or performance-dependent conditions.
Deciding on the degree of human control
Researchers need to balance the efficiency of AI with the ability to intervene and customize AI outputs. Researchers should decide how much flexibility they require in modifying AI outputs. For example, they might want to customize AI-generated codes, suggestions, or thematic summaries based on their interpretive insights. AI tools should offer options for manual adjustment and refinement to align with the researcher’s perspective.
AI integration across theoretical QDA frameworks
Our study suggests that researchers’ coding workflows may not strictly adhere to specific coding techniques or theoretical frameworks in everyday practice. Hence, designing AI support that is rigidly tied to a specific QDA approach may be counterproductive for many research projects. Instead, flexibility should be prioritized to accommodate diverse and evolving workflows. However, for research projects that rigorously follow a specific QDA framework, specialized or adjustable AI models could enhance research rigor and alignment with theoretical principles. Testing the feasibility and effectiveness of such AI systems could involve training models with data coded using a particular framework or instructing LLMs to adopt a designated theoretical approach when performing coding tasks. Furthermore, examining how AI-generated codebooks differ when different frameworks are applied may provide researchers with new insights into their data, supporting more informed decisions about the most suitable coding framework for their project. When AI is used as a supporting agent rather than the primary coder, its suggestions must align with the theoretical and methodological context defined by human coders. Misaligned suggestions can disrupt the research process and frustrate researchers by introducing inconsistencies. Ensuring compatibility between AI tools and the methodological requirements of a study is therefore crucial for maintaining coherence and usability in QDA workflows.
Evaluating Ethical Implications and Compliance
Integrating AI into QDA also involves evaluating ethical considerations, such as data privacy and security, to ensure compliance with ethical and legal standards and regulations. Additionally, participants in human subject studies should be informed during the consent process if AI will be used to analyze their data. Further, the harms of engaging with AI outputs need to be discussed, and safeguards should be applied when feeding data into AI systems. Finally, depending on the level of AI autonomy, a normative discussion on violations of intellectual integrity is necessary. For example, it may be necessary to report the level of AI usage in research outputs to maintain transparency and accountability in the research process. Researchers already use AI throughout their research process (e. g. (Kapania et al., 2024a; Argyle et al., 2023; Hämäläinen et al., 2023; Oppenlaender and Hämäläinen, 2023)), and popular QDA software increasingly integrates AI features – calling the scientific community, and specifically the CSCW and HCI community working at the intersection of humans and technology, to engage this debate and establish guidelines and best practices, e.g., on transparency. Kapania et al. (Kapania et al., 2024a) discuss the ethical guidelines for navigating these risks and concerns.
Human-Centered AI (HCAI) shares these ethical concerns and offers frameworks that prioritize human factors throughout the design, development, deployment, and use of AI systems (Usmani et al., 2023). Previous studies on automation highlight risks such as over-reliance and blind trust, often compounded by users’ reduced awareness and understanding of system behavior (“out-of-the-loop” effect (Xu et al., 2023)(Sarter and Woods, 1995; Xu, 2007). HCAI advocates human-in-the-loop AI systems and emphasizes “human-controlled autonomy”, advocating for AI systems that are designed to prioritize human oversight, transparency, and empowerment (Usmani et al., 2023; Xu et al., 2023). This approach aims to avoid excessive autonomy while addressing both the traditional challenges of automation and the novel complexities introduced by AI’s autonomous and adaptive characteristics (Shneiderman, 2020). By focusing on maintaining user control and awareness, HCAI promotes mechanisms for interpretable AI outputs, user feedback loops, and customizable settings to ensure that humans remain engaged and capable of intervention (Xu et al., 2023). These principles are highly relevant to the design of AI-assisted QDA workflows, aligning closely with our participants’ expressed concerns and conditions for AI involvement.

### v3

@@ 5.3. High AI Involvement: AI-Led Workflows
At this level, AI takes the lead in the QDA process, performing tasks autonomously with varying degrees of human oversight. However, participants raised ethical concerns, including the displacement of researchers and the risk of over-reliance on automated interpretation. These tensions underscore that full AI delegation remains contentious and highly context-dependent.
@@
In these configurations, AI performs core tasks such as initial coding or codebook generation, while researchers iteratively refine outputs through feedback. As P8 noted, “The AI suggests a codebook, and I give feedback that the tool uses to make an updated suggestion.” Cohn et al. (Cohn et al., 2024) propose a human-in-the-loop prompt engineering approach to collaborative discourse analysis. Dai et al. (Dai et al., 2023) develop an LLM-in-the-loop framework to develop an initial codebook.
AI with Human-in-the-Loop [Uncaptioned image] Figure 11. AI takes the lead in QDA, e. g., creating a codebook draft (left) or performing initial coding (right), with human oversight throughout. Figure 12 visualizes Role 12, AI with Human-in-the-Loop. It depicts 2 small flowcharts, visualizing 2 different possible workflows, each consisting of 3 elements. The first depicts pictogram of a computer, with an arrow leading to a cylinder including a tag, representing a codebook draft. An arrow leads to a pictogram of a person with a magnifier symbol and a pen symbol. From here, an arrow leads back to the computer pictogram. The second depicts pictogram of a computer, with an arrow leading to a pictogram of a paper sheet. A cylinder with an arrow and a tag leads to the paper sheet, representing initial coding. An arrow leads to a pictogram of a person with a magnifier symbol and a pen symbol. From here, an arrow leads back to the computer pictogram.
@@
Other participants envisioned workflows where AI acts autonomously but requires human approval, e.g., reviewing random subsamples. P13 suggested, “AI could apply my codebook, and I manually review a random subsample.” Recent studies use LLMs as additional coders that are evaluated in a final step (Tai et al., 2024; Bano et al., 2023; Xiao et al., 2023; Drápal et al., 2023).
AI with Human Approval [Uncaptioned image] Figure 12. AI creates a codebook draft (top) or assigns codes (bottom). A researcher approves before decisions are finalized and changes are applied. Figure 13 visualizes Role 13, AI with Human Approval. It depicts 2 small flowcharts, visualizing 2 different possible workflows, each consisting of 3 elements. The first depicts computer, with an arrow leading to a cylinder containing a tag, representing a codebook draft. From here, an arrow leads to a pictogram of a person with a paper sheet depicting 3 lines for text and 3 checkmarks, representing researcher approval. The second depicts a computer, with an arrow leading to a paper sheet with a cylinder and a tag next to it, representing code assignment. From here, an arrow leads to a pictogram of a person with a paper sheet depicting 3 lines for text and 3 checkmarks, representing researcher approval.
@@
AI with Conditional Autonomy [Uncaptioned image] Figure 13. AI autonomously handles routine coding but alerts researchers when it encounters data requiring nuanced understanding or deviating from the specified criteria. Figure 14 visualizes Task 14, AI with Conditional Autonomy. It depicts a small flowchart consisting of 3 elements. Element 1 depicts a computer. An arrow leads to element 2, a pictogram of a paper sheet with two tags and a question mark next to it. From the question mark, an arrow leads to element 3, a pictogram of a person.
Full or near-full AI Delegation
While no participant supported fully autonomous QDA end-to-end, some were open to delegating individual steps, such as applying a finalized codebook or generating an initial one (P8). Agentic AI systems and multi-agent approaches illustrate this direction (Rasheed et al., 2024). However, participants emphasized that such use cases demand heightened scrutiny due to risks around validity, reflexivity, and loss of human agency (Jiang et al., 2021; Feuston and Brubaker, 2021). OpenAI’s Viable can be used autonomously for analysis of unstructured data at scale, for example, in businesses that require regular customer feedback evaluations.
5.4. Summary and Implications
While HCI researchers were open to exploring the potential of AI support for QDA, they emphasized the need for responsible design that respects research ethics, quality, and researcher control. Our framework highlights that responsible AI integration in QDA is less about replacing human analysis and collaboration, and more about negotiating agency, trust, and epistemic responsibility. Notably, many participants articulated multiple ideas across different levels of integration, underscoring that acceptable AI involvement is fluid and context-dependent rather than fixed.
We stress that this framework is descriptive rather than normative: outlining how AI could be integrated does not imply how it should be used. Our positionality (early-career, WEIRD HCI researchers with experience in qualitative methods and AI) shapes this synthesis. Open questions remain about who benefits (or is harmed), what labor is displaced or amplified, and whether productivity should be the dominant metric guiding adoption. Our findings suggested that participants were open to it, and that a key driving factor was to increase productivity: raising difficult questions around whether speed should be be a primary goal in academic qualitative research.
At the same time, we do not view the framework as static. As new interaction paradigms emerge, such as more conversational, adaptive, or team-aware AI systems, and as collaborative research practices evolve, the roles and boundaries we identify will likely shift. We therefore expect the framework to be iteratively refined through future empirical work, serving as a living scaffold that evolves alongside changing human–AI relationships in qualitative research.
Overall, our findings underscore that AI support in QDA is neither neutral nor inevitable. Any integration must carefully weigh ethical, methodological, and practical considerations.
6. How our Framework Might be Used
Our framework aims to empower HCI and CSCW researchers to critically examine if, where, and how to incorporate AI tools into QDA. Rather than defaulting to AI-driven coding, it highlights multiple ways AI could support QDA while preserving research autonomy.
Determining the appropriate role of AI requires careful consideration of research goals, data types, and QDA methods. Below, we outline key considerations for using the framework to decide what role AI should play in a given study.
Determining the appropriate level of AI involvement
AI involvement in QDA is highly context-dependent and should thus not be treated as default. Researchers must first decide whether AI is appropriate at all, or whether other technologies could support the task. This requires literacy in both AI and QDA, particularly given risks of overreliance and explainability pitfalls (Feuston and Brubaker, 2021; Okolo et al., 2022; Buçinca et al., 2021). Such decisions should be made collaboratively within research teams. For example, large-scale or structured datasets may benefit from AI’s ability to handle pattern recognition, while nuanced or sensitive data might necessitate limited AI involvement to maintain interpretive depth and ethical considerations. Research goals – whether they focus on exploration, theory generation, or confirmation – also dictate how AI can best support QDA. AI may be more useful in early stages (e.g., surfacing patterns) or later stages (e.g., consistency checks), while in some cases it may best support technical tasks rather than analysis itself.
Deciding when AI should lead or support
Related work shows that coding performance can depend on various factors of complexity, e. g., the required level of judgment and interpretation  (Ortloff et al., 2023; Kirsten et al., 2024). Routine tasks such as transcription may warrant greater AI autonomy, whereas ambiguous or interpretive tasks call for stronger human leadership. In addition, researchers’ comfort and familiarity with AI tools as well as QDA should guide their level of involvement (Lubars and Tan, 2019; Gauthier and Wallace, 2022). Design strategies such as restricting the visibility of AI outputs or displaying performance disclaimers may further support appropriate use (Bhatt and Sargeant, 2024; Lam et al., 2024).
Identifying stages where human oversight is essential
Human oversight remains critical at key points in the workflow. Researchers should determine at which stages oversight is crucial -— such as after each coding pass, document review, or when salient patterns emerge. These checkpoints support collaborative sensemaking and ownership rather than replacing human analysis. When AI operates conditionally or autonomously, researchers should define clear triggers for human intervention to preserve agency.
Deciding on the degree of human control
Researchers must balance AI efficiency with the ability to intervene and customize outputs (Wang et al., 2020; Morris et al., 2024). AI tools should allow manual refinement of codes or summaries to align with researchers’ interpretive perspectives. However, excessive oversight can introduce new burdens (Feuston and Brubaker, 2021; Lubars and Tan, 2019; Gauthier and Wallace, 2022; Lam et al., 2024). In such cases, less AI involvement may be preferrable.
AI integration across theoretical QDA frameworks
Our findings suggest that researchers may not rigidly adhere to a single QDA framework in practice. AI systems tightly coupled to specific methods may therefore be counterproductive. Flexible tools that accommodate evolving workflows are likely more useful. However, projects grounded in specific frameworks may benefit from adjustable or specialized AI models, for example by instructing LLMs to adopt a given theoretical lens. Furthermore, examining how AI-generated codebooks differ when different frameworks are applied may provide researchers with new insights into their data, supporting more informed decisions about the most suitable coding framework for their project. When AI is used as a supporting agent rather than the primary coder, its suggestions must align with the theoretical and methodological context defined by human coders. Misaligned suggestions can disrupt the research process and frustrate researchers by introducing inconsistencies and workload. Ensuring compatibility between AI tools and the methodological requirements of a study is therefore crucial for maintaining coherence and usability in QDA workflows.
Evaluating Ethical Implications and Compliance
AI integration in QDA raises ethical concerns around privacy, security, consent, and intellectual ownership. Participants should be informed when AI is used in analysis, and safeguards must be in place when sharing data with AI systems. Finally, depending on the level of AI autonomy, a normative discussion on violations of intellectual integrity and ownership is necessary. For example, it is necessary to report the level of AI usage in research outputs to maintain transparency and accountability in the research process, especially in face of issues with transparency and reproducibility (Wacharamanotham et al., 2020; Fiesler et al., 2019).
Call to Action
Researchers already use AI throughout their research process (e. g. (Kapania et al., 2025; Argyle et al., 2023; Hämäläinen et al., 2023; Oppenlaender and Hämäläinen, 2023)), and popular QDA software increasingly integrates AI features. Our findings show that researchers do not seek a single best way to integrate AI into QDA. Instead, they emphasize choice, context, and control. This suggests that the key challenge for AI-assisted QDA is not increasing autonomy or speed, but supporting deliberate, situated decisions about whether and how AI should be involved at all.
@@ 7. Discussion
7.1. Synthesizing Prior Work on AI-Supported QDA
Our findings extend prior CSCW and HCI research on AI-supported Qualitative Data Analysis (QDA(Feuston and Brubaker, 2021; Jiang et al., 2021; Marathe and Toyama, 2018; Gao et al., 2023; Kapania et al., 2025; Zhang et al., 2024; Chatzichristos, 2025) by grounding this body of work in an empirical account of how researchers conduct QDA in practice. The diversity of QDA approaches, persistent skepticism toward AI integration (Jiang et al., 2021; Feuston and Brubaker, 2021; Chatzichristos, 2025), varied epistemic backgrounds of CSCW and HCI researchers, and the fundamentally social and collaborative nature of QDA (Fiesler et al., 2019; McDonald et al., 2019; Wallace et al., 2017; Gauthier and Wallace, 2022) underscore the need for careful and context-sensitive integration of AI into these workflows.
7.1.1. Workflows and Situated QDA Practice
While earlier work examined tensions around reliability, transparency, and reflexivity in qualitative research (Fiesler et al., 2019; McDonald et al., 2019), these conversations remain fragmented across subfields and methodological traditions. By documenting researchers’ workflows in detail, we synthesize these debates into an empirically grounded overview that makes visible recurring pain points – such as negotiating consensus or applying IRR – alongside the rewarding aspects of collaborative sensemaking. This workflow-centered account situates AI-supported QDA within CSCW’s long-standing recognition of qualitative analysis as a negotiated and social process (Wallace et al., 2017). Building on calls for flexible and customizable AI designs for QDA (Feuston and Brubaker, 2021), we extend prior work to the era of generative AI, highlighting emerging concerns such as hallucinations and shifting researcher expectations around AI assistance.
7.1.2. Shifting Attitudes toward AI Collaboration
Jiang et al. (2021) applied the AI task delegability framework (Lubars and Tan, 2019) to assess which QDA tasks AI should perform, considering factors like motivation, risk, trust, and difficulty. In line with research outside HCI and CSCW (Dellafiore et al., 2025), they found researchers to be largely cautious of AI intervention. In contrast, we observe growing openness to AI-based support, particularly when human agency and control are preserved. Recent work further identified a generational divide in openness, with early career researchers being more open, and more experienced researchers being more cautious (Chatzichristos, 2025). These patterns highlight the need for careful design that accommodates varying proficiencies in AI capability as well as QDA, reminding researchers on the limitations of AI, and allowing for evaluation and fine-tuning while avoiding excess of workload of doing so (Feuston and Brubaker, 2021; Lubars and Tan, 2019; Gauthier and Wallace, 2022; Lam et al., 2024). However, the field of Human-Centered Explainable AI is still in the making (Ehsan et al., 2024), and research shows deficiencies in meeting user needs and exacerbating issues of algorithmic opacity (Ehsan et al., 2021, 2022, 2023; Khalili, 2023). Okolo (Okolo et al., 2022) further found work on Explainable AI to be centered around the Global North, with concerning gaps in knowledge about the Global Souths, warning about “explainability pitfalls” (Ehsan and Riedl, 2024) causing overreliance on AI. Crucially, integrating AI into QDA introduces a shift from human-AI interaction to human-AI collaboration, that involves “mutual goal understanding, preemptive task co-management and shared progress tracking” (Wang et al., 2020) in sensemaking, and more research is needed how this can be best supported to avoid lopsided collaboration structures in AI-assisted QDA(Morris et al., 2024; Wang et al., 2020). Related work emphasizes this collaboration needs to be highly interactive, facilitated by mutual learning that fosters a symbiotic partnership that preserves researcher autonomy, rather than lending autonomy to AI (Feuston and Brubaker, 2021; Cho and Rader, 2020; Nagao, 2019). Our results have shown that collaborative AI roles were indeed perceived as favorable QDA scenarios by our study participants.
7.1.3. Messiness, Productivity, and the Limits of Automation
Our findings align with prior insights on the messiness and uncertainty of data analysis, which our participants also acknowledged. However, our interviews reveal that not all pain points are areas where researchers desire automation; for example, participants often prefer to engage directly with the “messiness” of their data to better understand it (Paulus and Marone, 2024; Jiang et al., 2021). Still, AI can support engagement with messiness at scale, bridging “small” and “big” data analysis approaches (Jiang et al., 2021; Lam et al., 2024; Baumer et al., 2017; Muller et al., 2016).
Consistent with Jiang et al. (2021), our participants emphasized the intrinsic value they find within the QDA process, including the satisfaction of uncovering patterns and learning through collaboration, which is particularly relevant in fields like CSCW and HCI which comprise researchers with various backgrounds and epistemic traditions (Jiang et al., 2021; McDonald et al., 2019; Bjørn and Boulus-Rødje, 2015; Fiesler et al., 2019). These insights challenge productivity-focused narratives surrounding AI, highlighting potential tensions between efficiency and QDA’s interpretive depth (Chatzichristos, 2025; Paulus and Marone, 2024), social dynamics, and opportunities for learning, skill-building, and knowledge transmission (Jiang et al., 2021).
7.2. Tensions, Values, and Inequities in AI-Supported QDA
While AI-supported QDA offers novel opportunities, concerns about output quality and bias remain central. These concerns are particularly salient when working with non-English data (Heseltine and Clemm von Hohenberg, 2024; Suter and Meckel, 2024), engaging minority communities, or conducting research in the Global Souths, as most AI systems are trained on English and other high-resource languages and are not “culturally agnostic” (Paulus and Marone, 2024). As a result, such systems often privilege dominant or high-frequency representations (Feuston and Brubaker, 2021), flatten linguistic nuance, and embed familiar Western styles of reasoning and analysis, making it harder to account for alternative narrative forms, cultural contexts, or theoretical traditions. Given that perspectives from the Global Souths and majority world are already marginalized within HCI and CSCW (Kumar and Karusala, 2021; Sturm et al., 2015; Linxen et al., 2021; Hasegawa et al., 2024), AI-supported QDA risks further amplifying these imbalances by embedding dominant interpretative frameworks and overlooking locally situated meanings (Kumar and Karusala, 2021). Thus, it is essential to build culturally sensitive and context-aware approaches to AI-supported QDA.
Finally, the integration of AI into QDA unfolds alongside broader debates in HCI and CSCW concerning citational justice (Kumar and Karusala, 2021), transparency of research practices (Wallace et al., 2017), and epistemic tensions (Kumar and Karusala, 2021; Fiesler et al., 2019; Pierre et al., 2021; Ajmani et al., 2024; Soden et al., 2024). Researchers caution against a “gradual return to a methodological positivism” (Chatzichristos, 2025) that AI could facilitate within interpretive QDA, while at the same time emphasizing the complementary potential of computational, positivist, and interpretive approaches (Baumer et al., 2017; Chatzichristos, 2025; Gauthier and Wallace, 2022; Lam et al., 2024; Muller et al., 2016). Situated within broader politics of knowledge production in HCI and CSCW (Kumar and Karusala, 2021; Collective et al., 2021; Soden et al., 2024), these dynamics are further shaped by uneven access to AI-enabled QDA tools, which often require reliable internet access, computational resources, or paid subscriptions; advantages that disproportionately accrue to well-resourced institutions and researchers. These tensions call for sustained reflection on epistemic justice and geographically shaped access to resources (Chatzichristos, 2025; Ajmani et al., 2024; Kumar and Karusala, 2021), including tools, time, and funding, to avoid reinforcing existing inequalities through AI integration into QDA.